# A Survey on Dynamic Neural Networks: from Computer Vision to Multi-modal Sensor Fusion

### v3

@@ 4 Discussion
Since the rise of deep learning solutions in Computer Vision, Dynamic Neural Network models have seen increasing adoption, as highlighted by the increasing trend of publications over time (cf. Figure 2). The topic of Early Exits is the most popular within the relevant literature, with a large portion of the research effort spent in understanding both the different types of architectures and the different improvement aspects of those: from optimal exiting policies and positioning, correct weighting for the loss contribution of intermediate classifiers, mismatches between the training and inference phase, and so on. Based on the analysis conducted in this survey, the number of publications in this area has consistently grown since 2016, with activity stabilizing after 2020.
For techniques involving dynamic computational paths, the research landscape is extensive but more diverse in its approaches. Different types of solutions have been presented: from networks that are constructed on the fly by composing different computation blocks according to the input, to the dynamic skip of certain computation units, to the sub-specialization of specialized “expert” blocks activated only to solve certain situations, in the case of MoE. The topic has seen a peak in publications around 2018, targeting more the dinamicity of CNNs blocks, followed by a peak in 2020, with the rise of MoE techniques aimed at enabling efficient training of larger models.
@@
Challenges and limitations
@@
Benchmarks and metrics. The lack of standardized metrics and benchmarks tailored to Dynamic Neural Networks complicates the comparison among different methods. Conventional metrics (e.g., accuracy, FLOPs) are often not sufficient to capture the adaptive nature of these models. Currently, the literature relies on comparisons within the most similar architecture, making a broader comparison among different approaches hard to make.
@@ 5 Conclusions
This survey provides a comprehensive analysis of 161 Dynamic Neural Network methodologies for Computer Vision and their applications in Sensor Fusion. We propose a taxonomy categorizing the techniques according to their adaptive network component: Early Exits (output-level), Dynamic Routing (path-level), and Token Skimming (input-level). Furthermore, we examine the specific advantages of these architectures for multi-modal tasks and supplement this review with a comprehensive, up-to-date repository of implementations to facilitate future research. All models that constrained the network output position with respect to the input complexity fall under the Early Exits Section, where we presented the most important architectural solutions, relevant methods for improving the training, and the application of these methods to specific tasks and use cases. Then we proceeded with the review of Dynamic Routing techniques, where the computational path is created dynamically, either by constructing it on the fly or routing the computation to different model blocks. Also, in this case, a we present the application of these methods to various tasks. Lastly, in the Token Skimming section, we focused on Transformer-based models that either drop certain tokens along the block of the network (temporarily or permanently) or merge them together, based on an estimated token contribution to the desired task. The last part is dedicated to other methods and the applications of skimming to tasks that differ from image classification.
@@
Furthermore, we provided a discussion on the current limitations and future directions. A considerable highlight is that Token Skimming allows for direct comparison on the ImageNet benchmark, where SaiT currently holds the best overall accuracy, whereas Early Exits and Dynamic Routing still lack a shared standard for consistent evaluation.

### v4

@@ 4 Discussion
Since the rise of deep learning solutions in Computer Vision, Dynamic Neural Network models have seen increasing adoption, as highlighted by the increasing trend of publications over time (cf. Figure 2). The topic of Early Exits is the most popular within the relevant literature, with a large portion of the research effort is spent in understanding both the different types of architectures and the different improvement aspects of those: from optimal exiting policies and positioning, correct weighting for the loss contribution of intermediate classifiers, mismatches between the training and inference phase, and so on. Based on the analysis conducted in this survey, the number of publications in this area has consistently grown since 2016, with activity stabilizing after 2020.
For techniques involving dynamic computational paths, the research landscape is extensive but more diverse in its approaches. Different types of solutions have been presented: from networks that are constructed on the fly by composing different computation blocks according to the input, to the dynamic skip of certain computation units, to the sub-specialization of specialized “expert” blocks activated only to solve certain situations, in the case of MoE. The topic has seen a peak in publications around 2018, targeting more the dynamicity of CNN blocks, followed by a peak in 2020, with the rise of MoE techniques aimed at enabling efficient training of larger models.
@@
4.1 Standardized evaluation template for Dynamic Neural Networks
As discussed throughout this survey, the domain of Dynamic Neural Networks currently lacks a standardized experimental protocol. The current literature largely relies on comparisons within the same category of methodologies and often utilizing a diverse array of datasets and evaluation baselines which makes comparisons unfeasible. This directly translates into infeasible quantitative analysis without extensive re-implementation. Furthermore, conventional metrics, such as static accuracy and overall FLOPs, are often not sufficient to capture the adaptive nature of these models. We argue that a structured benchmark is essential to consistently compare different Dynamic Neural Network models. Such an evaluation must specifically account for actual inference speed and efficiency, the variability of computations in relation to input complexity, and performance quality in edge cases. To address this critical gap and facilitate future cross-study analysis, we propose a draft formulation of a standardized evaluation protocol for reporting quantitative results. We believe that there are four core aspects that a standardized template must address: the trade-off between performance and efficiency, theoretical versus hardware efficiency, a quantification of the dynamicity and a direct comparison with a static equivalent baseline. We proceed by detailing these proposed criteria below.
Performance-efficiency trade-off
Static networks yield a single operational point in terms of accuracy and computational cost (FLOPs). In contrast, Dynamic Neural Networks inherently yield a continuous trade-off curve based on their budget thresholds (e.g., confidence levels in Early Exits). To establish a benchmark that is strictly table-comparable without relying on arbitrary discrete ”cherry picked” configurations, we propose standardizing evaluations around two rigorous metrics: the Area Under the Curve (AUC) of the accuracy-computation trade-off, and the computational variance. Utilizing the AUC summarizes the network’s overall resilience across all computational budgets into a single metric. Reporting the average computational cost (e.g., Average GFLOPs) alongside its standard deviation allows to quantify the network’s dynamicity, proving that it actively redistributes computations based on sample complexity.
Computational efficiency versus hardware reality
There is a known discrepancy in the current literature between theoretical computational cost and actual inference speed. A dynamic router might theoretically save 30% of FLOPs, but inadvertently slow down the actual inference due to fragmented memory access. Hardware accelerators, such as GPUs and TPUs, inherently favor static, dense computations. Furthermore, memory bandwidth and data movement costs often dominate the actual inference latency in edge devices. We argue benchmarks must explicitly state the hardware platform utilized and measure actual wall-clock time. To further standardize these empirical metrics, we encourage conducting evaluations on integrated System-on-Chip (SoC) platforms (e.g., the NVIDIA Jetson family). The use of these integrated platforms allows to reduce the measurement noise generated by the different hardware components (e.g. CPU, GPU, memory) present in standard desktop setups. Furthermore, when evaluating models intended for embedded systems, the energy consumption should also be reported.
Quantifying dynamicity
A critical, yet often overlooked, aspect of evaluation is the actual dynamic behavior of the network. A network is dynamic if the computational path can be adjusted to the input complexity. To effectively quantify this adaptability, we propose the introduction of specific metrics evaluated on a fixed, representative network configuration (e.g., a default or optimal budget threshold). Specifically, we suggest reporting the compute variance across the test set. Evaluating a single operational point allows to isolate and measure the true adaptivity of the model to the complexity of individual samples, rather than the variance generated by different hyperparameter settings.
Fair baselines
When evaluating Dynamic Neural Networks, a direct comparison with their static counterpart enables a better quantification of the type and magnitude of the improvement introduced by the dynamic components. The direct comparison offers a clear intuition of the structural and computational advantages provided by the dynamicity with respect to a well-known architecture. Furthermore, the static baseline can serve as an effective proxy to compare different dynamic methodologies. In a heterogeneous literature landscape where training setups frequently vary, evaluating the relative improvements of different dynamic networks over the exact same static backbone (e.g., analyzing two distinct dynamic routing methods that both build upon a standard ResNet-50) allows to contextualize the results and provides a fairer ground for cross-paper comparisons.
Table 6: A proposed standardized evaluation template applied to landmark Dynamic Neural Networks. Extracting these metrics highlights historical inconsistencies: while recent Vision Transformer methods correctly report empirical Image Throughput, earlier literature often poorly standardized latency, and crucially, none report AUC, computational variance, or energy metrics. Standardizing these fields is essential for fair, comprehensive cross-family comparisons. Abbreviations: $E$: Early Exits, DR: Dynamic Routing, TD: Token Drop, TM: Token Merge. RN: ResNet, DN: DenseNet. C-10: CIFAR-10, IN-1K: ImageNet-1K. GF: GFLOPs. $\sigma$: Standard Deviation. AUC: Area Under the Curve.
Method Type Arch. Data Hardware Static (Acc / GF) Dynamic (Acc / Avg GF $\pm\sigma$) AUC (Acc-GF) Comput. Savings Inference Speed Energy (mJ/im) BranchyNet EE RN-110 C-10 Titan X 80.7% / – – / – ($\pm$ –) 37.2 ms MSDNet EE Custom IN-1K $\sim$75.0% / 1.7 ($\pm$ –) DVT EE T2T-ViT-12 IN-1K 2080 Ti 76.7% / 1.78 76.2% / 1.00 ($\pm$ –) $\sim$43% 1128 im/s SkipNet DR RN-101 IN-1K 77.4% / – – / – ($\pm$ –) 30% BlockDrop DR RN-101 IN-1K P6000 76.4% / 15.6 76.4% / 12.5 ($\pm$ 0.43) 19.9% ConvNet-AIG DR RN-101 IN-1K 76.4% / 7.6 77.5% / 5.1 ($\pm$ –) 38% DynamicViT TD LV-ViT-S IN-1K RTX 3090 83.3% / 6.6 83.0% / 4.6 ($\pm$ –) 31% 1417 im/s EVIT TM DeiT-S IN-1K A100 79.8% / 4.6 79.5% / 3.0 ($\pm$ 0.0) 35% 4385 im/s ToMe TM ViT-B (Aug) IN-1K V100 84.6% / – 83.9% / – ($\pm$ 0.0) 406 im/s
To demonstrate the usefuleness of the proposed framework, Table 6 presents a retrospective application of our criteria to a selection of landmark Dynamic Neural Networks. For this analysis, we selected three of the most representative publications for each main methodological category. Extracting these metrics directly from the original manuscripts highlights several inconsistencies and gaps in standard evaluation protocols. As evident from the missing values, none of the considered foundational works report the Area Under the Curve (AUC) or energy consumption, and only a minority explicitly detail the computational variance ($\sigma$). Furthermore, the table illustrates a clear evolution in how empirical inference speed is quantified. In earlier literature, such as ResNet-era routing architectures, GFLOPs and latency (e.g., in milliseconds) were often poorly standardized or omitted entirely. In contrast, more recently image throughput (images/sec) has become the standardized metric over single-image latency. Our proposed evaluation template highlights what should be reported to bridge reporting gaps, ensuring that future architectures are benchmarked comprehensively across theoretical sparsity, dynamic adaptivity, and practical hardware utility.
4.2 Other challenges and limitations
@@ 5 Conclusions
This survey provides a comprehensive analysis of 163 Dynamic Neural Network methodologies for Computer Vision and their applications in Sensor Fusion. We propose a taxonomy categorizing the techniques according to their adaptive network component: Early Exits (output-level), Dynamic Routing (path-level), and Token Skimming (input-level). Furthermore, we examine the specific advantages of these architectures for multi-modal tasks and supplement this review with a comprehensive, up-to-date repository of implementations to facilitate future research. All models that constrained the network output position with respect to the input complexity fall under the Early Exits Section, where we presented the most important architectural solutions, relevant methods for improving the training, and the application of these methods to specific tasks and use cases. Then we proceeded with the review of Dynamic Routing techniques, where the computational path is created dynamically, either by constructing it on the fly or routing the computation to different model blocks. Also, in this case, a we present the application of these methods to various tasks. Lastly, in the Token Skimming section, we focused on Transformer-based models that either drop certain tokens along the block of the network (temporarily or permanently) or merge them together, based on an estimated token contribution to the desired task. The last part is dedicated to other methods and the applications of skimming to tasks that differ from image classification.
@@
Furthermore, we provided a discussion on the current limitations and future directions. A considerable highlight is that Token Skimming allows for direct comparison on the ImageNet benchmark, where SaiT currently holds the best overall accuracy, whereas Early Exits and Dynamic Routing still lack a shared standard for consistent evaluation. For other method, a straight comparison is not trivial. We formulate a standardized evaluation template to address the lack of unified benchmarking protocols, with the scope of providing a fair and rigorous cross-methodology comparison methodology for future publications.